# A solution to Banach conjecture

### v8

@@ 2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Gardner [6] and Schneider [14].
@@ 3. Proof of the main theorem
$\begin{cases}E_{\alpha}:=\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cap\alpha(K\cap\xi_{0}^{\perp})\big),\\[4.0pt] \tilde{E}_{\beta}:=\displaystyle\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup\beta(K\cap\xi_{0}^{\perp})\big).\end{cases}$
@@
respectively, here $\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi^{\perp}$. Therefore, to prove that $E_{\alpha}$ and $\tilde{E}_{\beta}$ are continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
@@
Remark 3.2.
Observe that the continuous map $\Phi:\mathbb{S}^{n-1}\to GL(n)/G_{K\cap\xi_{0}^{\perp}}$ admits a continuous lift $\phi:\mathbb{S}^{n-1}\to GL(n)$ such that $\phi(\xi_{0})$ is the identity and, for every $\xi$, we can write $\phi(\xi)=\phi_{\xi}\circ g_{\xi}$ with $g_{\xi}\in G_{K\cap\xi_{0}^{\perp}}$. Moreover, both $K\cap\xi_{0}^{\perp}$ and $J(K\cap\xi_{0}^{\perp})$ are preserved by $g_{\xi}$. Hence, without loss of generality, we will from now on take $\phi_{\xi}$ to be this lift $\phi(\xi)$.
@@
$\Gamma:=\big\{(\theta,\xi,t)\in\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]:\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}=t\ \text{and}\ \theta\perp\xi\big\}.$
Observe that, for any $\xi\in\mathbb{S}^{n-1}$, Lemma 3.1 implies
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}$
$=$ $\big\|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\theta\big\|_{K\cap\xi^{\perp}}$
$=$ $\big\|\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\theta:\lambda>0\}\big\|_{K\cap\xi^{\perp}}$
$=$ $\big\|J(K\cap\xi_{0}^{\perp})\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{\phi^{-1}_{\xi}(K\cap\xi^{\perp})}$
$=$ $\big\|J(K\cap\xi_{0}^{\perp})\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{K\cap\xi_{0}^{\perp}}$
$=$ $\rho_{J(K\cap\xi_{0}^{\perp})}(\phi^{-1}_{\xi}(\theta))\big\|\phi^{-1}_{\xi}(\theta)\big\|_{K}.$
And hence, for any $\eta\in\mathbb{S}^{n-1}$
$(\theta,\xi,t)\in\Gamma\iff(\phi^{-1}_{\xi}(\theta),\xi_{0},t)\in\Gamma\iff(\phi_{\eta}\circ\phi^{-1}_{\xi}(\theta),\eta,t)\in\Gamma$
From this point of view, we can construct a continuous mapping $\Lambda_{t}[\theta]:\mathbb{S}^{n-1}\to\mathbb{S}^{n-1}$, defined by
$\Lambda_{t}[\theta](\eta)=\phi_{\eta}\phi^{-1}_{\xi}(\theta).$ (1)
@@
$\Lambda_{t}:=\{\theta\in\mathbb{S}^{n-1}:(\theta,\xi,t)\in\Gamma\mbox{ for some }\xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\}.$ (2)
It is clear that $\Lambda_{t}[\theta](\eta)\subseteq\Lambda_{t}$ for any $\theta\in\Lambda_{t}$. We now present several basic propositions of $\Lambda_{t}[\theta]$ and $\Lambda_{t}$.
@@
    Either $\Lambda_{t}$ is the entire sphere, or the image of $\Lambda_{t}[\theta]$ is null-homotopic and its complement $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ consists of exactly two connected components;
@@
    Either $\Lambda_{t}$ is the entire sphere, or it is null-homotopic, symmetric with respect to the origin, and contains an equator;
@@
Before presenting the proof, we note that the $\phi_{\xi}$ appearing here is the continuous lifting introduced in Remark 3.2. After this proposition, however, the continuity of $\phi_{\xi}$ is no longer required, since the set $\Lambda_{t}$ does not depend on the continuity of $\phi_{\xi}$.
@@
First, we aim to prove that if $\Lambda_{t}$ is a proper subset of the sphere, then $\Lambda_{t}[\theta]$ is null-homotopic. In fact, if $a\notin\Lambda_{t}$, then $a\notin\Lambda_{t}$ also holds because $K$ is symmetric about the origin. Thus, $a,-a\notin\mbox{Im}(\Lambda_{t}[\theta])$, and hence we can define a homotopy
$F(\eta,s)=\frac{s\,\Lambda_{t}[\theta](\eta)+(1-s)a}{\|s\,\Lambda_{t}[\theta](\eta)+(1-s)a\|}.$
@@
Observe that for any $\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)\in\Lambda_{t}[\theta]$ and any $\zeta\in\mathbb{S}^{n-1}\cap(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta))^{\perp}$, the set
$\{\phi_{\vartheta}\circ\phi^{-1}_{\eta}(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)):\vartheta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta),\zeta)\}\subseteq\mathrm{Im}(\Lambda_{t}[\theta])$
traces out a curve that passes through $\phi_{\eta}\circ\phi_{\xi}^{-1}(\theta)$ and enters both $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{+}$ and $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{-}$ (see Figure 1).
Figure 1. The level set $\Lambda_{t}[\theta]$ passes through $\zeta^{\perp}$.
Consequently, every boundary point $x$ of one branch $C$ of $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ must lie on a geodesic ball that is tangent at $x$ and has radius $r_{x}<\frac{\pi}{2}$. This subsphere can be written as
@@
Otherwise, we may prolong a curve contained in $\mathrm{Im}(\Lambda_{t}[\theta])$ until it intersects $C$. By the compactness of $\bar{C}$, this yields a uniform radius $r_{0}<\frac{\pi}{2}$ such that $C$ lies entirely in a geodesic ball of radius $r_{0}$. Note that $\mathrm{Im}(\Lambda_{t}[\theta])$ is null-homotopic and $a,-a\notin\mathrm{Im}(\Lambda_{t}[\theta])$. Consequently, the complement $\mathbb{S}^{n-1}\setminus\mathrm{Im}(\Lambda_{t}[\theta])$ can only decompose into two connected components, $C_{1}$ and $C_{2}$, which contain $a$ and a-a, respectively. Now we prove Proposition (1).
For (2), assume $\Lambda_{t}$ is not the entire sphere. Then for any $\theta\in\Lambda_{t}$, Proposition (1) tells us that $\mathrm{Im}(\Lambda_{t}[\theta])$ is null-homotopic. Consequently, for any $\theta_{1},\theta_{2}\in\Lambda_{t}$, we must have
$\mathrm{Im}(\Lambda_{t}[\theta_{1}])\cap\mathrm{Im}(\Lambda_{t}[\theta_{2}])\neq\emptyset;$
@@
$\mathrm{Im}(\Lambda_{t}[\theta_{2}])\subseteq C_{1},$
[…]
@@
for some $r_{2}<\frac{\pi}{2}$. Therefore, $\theta_{1}$ and $\theta_{2}$ are path-connected within $\Lambda_{t}$. Let $\gamma$ be a path in $\Lambda_{t}$ joining them. Then
$\bigcup_{\theta\in\gamma}\mathrm{Im}(\Lambda_{t}[\theta])$
itself has to be null-homotopic. Moreover, for any $(\theta,\xi,t)\in\Gamma$, the symmetry of $K$ implies that $(-\theta,-\xi,t)\in\Gamma$ as well, which shows that $\Lambda_{t}$ is origin-symmetric. Therefore, $\mathbb{S}^{n-1}\setminus\Lambda_{t}$ consists of two antipodal connected components, each of which lies inside one of the geodesic balls
@@
for some $r<\frac{\pi}{2}$, respectively. Consequently, $\Lambda_{t}$ contains the equator $\mathbb{S}^{n-1}\cap o^{\perp}$. This establishes Proposition (2).
For (3), observe that there must exist some $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$ with $(\theta,\xi_{0},1)\in\Gamma$. If this were not the case, then $J(K\cap\xi_{0}^{\perp})$ would lie strictly inside $K\cap\xi_{0}^{\perp}$, contradicting the fact that $J(K\cap\xi_{0}^{\perp})$ is the John ellipsoid. If $\Lambda_{1}$ is not equal to the entire sphere, then $\Lambda_{1}$ is null-homotopic and contains some equator. For sufficiently small $\epsilon>0$, the union $\bigcup_{t\in(1-\epsilon,1]}\Lambda_{t}$ forms an open set $\mathcal{O}$ that contains $\Lambda_{1}$. Assuming that none of the $\{\Lambda_{t}\}_{t\in[1-\epsilon,1]}$ equals the whole sphere, we assert that
$\Lambda_{1}\subseteq\Lambda_{1-\epsilon}.$
If this were false, then $\Lambda_{1-\epsilon}$ would provide a null-homotopy that contains $\partial\mathcal{O}$, with $\Lambda_{1}\subseteq\mathcal{O}$ but omitting some points of $\Lambda_{1}$, which is a contradiction.
Likewise, for any $\epsilon>0$, assuming that none of the sets in the family $\{\Lambda_{t}\}_{t\in[1-\epsilon,1]}$ coincides with the entire sphere, it follows that
$\Lambda_{q}\subseteq\Lambda_{p}$
whenever $1-\epsilon\leq p\leq q\leq 1$.
We define
@@
Then any $t$ for which $\Gamma$ holds must lie in the interval $[m,1]$. Indeed, if this were not the case, we could find $(\theta,\xi,t)\in\Gamma$ with $t<m$, which would imply $(\phi^{-1}_{\xi}(\theta),\xi_{0},t)\in\Gamma$, contradicting the fact that $m$ is the infimum. Thus, we obtain
$\bigcup_{t\in[m,1]}\Lambda_{t}=\mathbb{S}^{n-1}.$
Consequently, there must exist some $t_{0}$ such that
$\Lambda_{t_{0}}=\mathbb{S}^{n-1}.$
Indeed, if this were not the case, then each $\Lambda_{t}$ for $t\in[m,1]$ would be a proper subset of the sphere. However, by the preceding argument,
$\Lambda_{m}=\bigcup_{t\in[m,1]}\Lambda_{t}=\mathbb{S}^{n-1},$
which is a contradiction. ∎
@@
$=\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)$
@@
[…]

### v9

@@ 2. Notation and Preliminaries
This section mainly presents some basic content, such as symbols and definitions. For more detailed content, readers can refer to the books by Fuente [6], Gardner [7], and Schneider [14].
@@ 3. Proof of the main theorem
$\begin{cases}E_{\alpha}:=\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cap\alpha(K\cap\xi_{0}^{\perp})\big),\\[4.0pt] \tilde{E}_{\beta}:=\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup\beta(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}.\end{cases}$
@@
respectively, here $\phi_{\xi}(K\cap\xi_{0}^{\perp})=K\cap\xi^{\perp}$. Therefore, to prove that the radial functions of $E_{\alpha}$ and $\tilde{E}_{\beta}$ are continuous, it suffices to establish the continuity of the mapping $\xi\mapsto\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))$.
@@
For a fixed $\theta\in\mathbb{S}^{n-1}\cap\xi^{\perp}_{0}$, we introduce the set $\tilde{\Lambda}[\theta]\subseteq\mathbb{S}^{n-1}$ by
$\tilde{\Lambda}[\theta]:=\{\,\phi_{\xi}(\theta):\phi_{\xi}\in\Phi(\xi)\ \text{and}\ \xi\in\mathbb{S}^{n-1}\,\}.$
Observe that, for any $\xi\in\mathbb{S}^{n-1}$, Lemma 3.1 implies
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}$
$=$ $\big\|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\eta\big\|_{K\cap\xi^{\perp}}$
$=$ $\big\|\partial(\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))\cap\{\lambda\eta:\lambda>0\}\big\|_{K\cap\xi^{\perp}}$
$=$ $\big\|\partial(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\phi^{-1}_{\xi}(\theta):\lambda>0\}\big\|_{\phi^{-1}_{\xi}(K\cap\xi^{\perp})}$
$=$ $\big\|\partial(J(K\cap\xi_{0}^{\perp}))\cap\{\lambda\phi^{-1}_{\xi}(\eta):\lambda>0\}\big\|_{K\cap\xi_{0}^{\perp}}$
$=$ $\rho_{J(K\cap\xi_{0}^{\perp})}(\phi^{-1}_{\xi}(\eta))\big\|\phi^{-1}_{\xi}(\eta)\big\|_{K}.$
And hence, for any $\eta\in\tilde{\Lambda}[\theta]$, there exists a $\xi\in\mathbb{S}^{n-1}\cap\eta^{\perp}$ such that
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}$
@@
$\Gamma:=\big\{(\eta,\xi,t)\in\mathbb{S}^{n-1}\times\mathbb{S}^{n-1}\times[0,1]:\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\eta)\|\eta\|_{K}=t\ \text{and}\ \eta\perp\xi\big\}.$
@@
$\Lambda_{t}:=\{\eta\in\mathbb{S}^{n-1}:(\eta,\xi,t)\in\Gamma\mbox{ for some }\xi\in\mathbb{S}^{n-1}\cap\eta^{\perp}\}.$ (1)
It is clear that $\tilde{\Lambda}[\theta]\subseteq\Lambda_{t}$ if $\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}=t$.
Observe that the linear symmetry group is compact ([3, Lemma 2.1]), and therefore has only finitely many connected components. We select $H\subseteq G_{K\cap\xi_{0}^{\perp}}$ to be the connected component containing the identity element. By the continuity of $\Phi$, for each $\eta\in\Lambda[\theta]$ with $\phi_{\xi}\circ g_{\xi}(\theta)=\eta$, we can choose a local continuous map $\phi_{\varphi}\circ g_{\varphi}:U_{\xi}\to GL(n)$, where $U_{\xi}$ is a neighborhood of $\xi$. Fixing $\phi_{\xi_{0}}$ to be the identity, we can construct a continuous path $\gamma$ from $\theta$ to $\eta$ by piecing together these local continuous maps. This path yields
$\phi^{-1}_{\xi}(\eta)\subseteq\{g(\theta):g\in H\}.$
We redefine
$\Lambda_{t}[\theta]:=\{$ $\phi_{\xi}(\zeta):\phi_{\xi}\in\Phi(\xi)\mbox{ is local continuous},\ \zeta\in\{g(\theta):g\in H\},$
$\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}=t,\ \text{and}\ \xi\in\mathbb{S}^{n-1}\}.$
We now present several basic propositions of $\Lambda_{t}[\theta]$ and $\Lambda_{t}$.
@@
    Either $\Lambda_{t}$ is the entire sphere, or the set of $\Lambda_{t}[\theta]$ is null-homotopic and its complement $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ consists of exactly two connected components;
@@
    Either $\Lambda_{t}$ is the entire sphere, or $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ is null-homotopic, symmetric with respect to the origin, and contains an equator;
@@
First, we aim to prove that if $\Lambda_{t}$ is a proper subset of the sphere, then $\Lambda_{t}[\theta]$ is null-homotopic. In fact, if $a\notin\Lambda_{t}$, then $a\notin\Lambda_{t}$ also holds because $K$ is symmetric about the origin. Thus, $a,-a\notin\Lambda_{t}[\theta]$, and hence we can define a homotopy
$F(\xi,g,s)=\frac{s\,\phi_{\xi}\circ g(\theta)+(1-s)a}{\|s\,\phi_{\xi}\circ g(\theta)+(1-s)a\|}.$
@@
Observe that for any $\eta\in\Lambda_{t}[\theta]$ and any $\zeta\in\mathbb{S}^{n-1}\cap\eta^{\perp}$, the set
$\{\phi_{\vartheta}\circ g_{\vartheta}(\theta)):\vartheta\in\mathbb{S}^{n-1}\cap\mathrm{span}(\eta,\zeta)\}\subseteq\Lambda_{t}[\theta]$
traces out a curve that passes through $\eta$ and enters both $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{+}$ and $\mathbb{S}^{n-1}\cap\zeta^{\perp}_{-}$ (see Figure 1).
Figure 1. The set $\Lambda_{t}[\theta]$ passes through $\zeta^{\perp}$.
Consequently, every boundary point $x$ of one branch $C$ of $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ must lie on a geodesic ball that is tangent at $x$ and has radius $r_{x}<\frac{\pi}{2}$. This subsphere can be written as
@@
Otherwise, we may prolong a curve contained in $\Lambda_{t}[\theta]$ until it intersects $C$. By the compactness of $\bar{C}$, this yields a uniform radius $r_{0}<\frac{\pi}{2}$ such that $C$ lies entirely in a geodesic ball of radius $r_{0}$. Note that $\Lambda_{t}[\theta]$ is null-homotopic and $a,-a\notin\Lambda_{t}[\theta]$. Consequently, the complement $\mathbb{S}^{n-1}\setminus\Lambda_{t}[\theta]$ can only decompose into two connected components, $C_{1}$ and $C_{2}$, which contain $a$ and a-a, respectively. Now we prove Proposition (1).
For (2), assume $\Lambda_{t}$ is not the entire sphere. Then for any $\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}$, we have $\Lambda_{t}[\theta]=-\Lambda_{t}[-\theta]$ by the symmetry of $K$. Proposition (1) also tells us that $\Lambda_{t}[\theta]$ and $\Lambda_{t}[-\theta]$ are null-homotopic. Consequently, we must have
$\Lambda_{t}[\theta]\cap\Lambda_{t}[-\theta]\neq\emptyset;$
@@
$\Lambda_{t}[-\theta]\subseteq C_{1},$
where $C_{1}$ denotes a connected component of $\mathbb{S}^{n-1}\backslash\Lambda_{t}[\theta]$. In this situation, there exists a connected component $\tilde{C}_{1}$ of $\mathbb{S}^{n-1}\backslash\Lambda_{t}[-\theta]$ that contains $\mathbb{S}^{n-1}\backslash C_{1}$; hence $\tilde{C}_{1}$ includes a large geodesic ball of the form
@@
for some $r_{2}<\frac{\pi}{2}$. Therefore, both $\Lambda_{t}[\theta]$ and $\Lambda_{t}[-\theta]$ are null-homotopic with $\Lambda_{t}[\theta]\cap\Lambda_{t}[-\theta]\neq\emptyset$ and $\Lambda_{t}[\theta]=-\Lambda_{t}[-\theta]$. And hence, $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ is null-homotopy and origin-symmetric. Therefore, $\mathbb{S}^{n-1}\setminus(\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta])$ consists of two antipodal connected components, each of which lies inside one of the geodesic balls
@@
for some $r<\frac{\pi}{2}$, respectively. Consequently, $\Lambda_{t}[\theta]\cup\Lambda_{t}[-\theta]$ contains the equator $\mathbb{S}^{n-1}\cap o^{\perp}$. This establishes Proposition (2).
For (3), we consider the continuous function
$h(\theta)=\rho_{J(K\cap\xi_{0}^{\perp})}(\theta)\|\theta\|_{K}.$
It is evident that $h$ is Lipschitz since both $K\cap\xi_{0}^{\perp}$ and $J(K\cap\xi_{0}^{\perp})$ are convex bodies in $\xi_{0}^{\perp}$.
We now assume the opposite: that $\Lambda_{t}\neq\mathbb{S}^{n-1}$ for any $t\in[m,1]$, where
@@
Note that for any connected level set of $h$, denoted by $\gamma$, if $\theta\in\gamma$, then
$\{g(\theta):g\in H\}\subseteq\gamma.$
Now we choose a connected level set of $h$, denoted by $\gamma$, that divides the sphere into finitely many connected components. If, instead, there were infinitely many connected components, then the compactness of $\mathbb{S}^{n-1}$ implies that some of them would have to be arbitrarily small, which would contradict the requirement that each component is an open set.
Thus, we may assert that there exists at least one connected component $C$ of $\mathbb{S}^{n-1}\setminus\gamma$ such that
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}\cup-\overline{C}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
Indeed, if this were not the case, then by Proposition (2), the condition $\cup\{\Lambda_{t}[\theta]:\theta\in\bar{C}\cup-\bar{C}\ \text{and}\ t\in[m,1]\}\neq\mathbb{S}^{n-1}$ would imply that
$\bigcup\{\Lambda_{t}[\theta]:\theta\in C\cup-C\ \text{and}\ t\in[m,1]\}$
is an open, null-homotopic subset of $\mathbb{S}^{n-1}$ that contains some equator, whose boundary is contained in
$\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}$
for some $\tilde{t}\in[m,1]$. Consequently, by Proposition (2), we must have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in C\cup-C\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\};$
otherwise, the set $\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}$ would fail to be null-homotopic. This shows that, for any connected component $C$ of $\mathbb{S}^{n-1}\setminus\gamma$, we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\bar{C}\cup-\bar{C}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\},$
and therefore
$\mathbb{S}^{n-1}=\bigcup\{\Lambda_{t}[\theta]:\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{\tilde{t}}[\theta]:\theta\in\gamma\cup-\gamma\}\subseteq\Lambda_{\tilde{t}}$,
which contradicts our assumption.
We now consider the component $C$ for which
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}\cup-\overline{C}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
Then We introduce a partially ordered set $(\mathcal{G},\prec)$ with the following structure: for each $\alpha\in\mathcal{G}$, the curve $\gamma_{\alpha}$ is a connected level set of $\{\theta\in\mathbb{S}^{n-1}\cap\xi_{0}^{\perp}:h(\theta)=t_{\alpha}\}$, and $C_{\alpha}$ denotes one of the connected components of the complement $C\setminus\gamma_{\alpha}$ whose boundary lies on $\gamma_{\alpha}$. We then declare $\alpha\prec\beta$ whenever $C_{\alpha}\subsetneq C_{\beta}$.
observe that if
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1},$
for any $\beta\succ\alpha$, we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}=\mathbb{S}^{n-1}.$
On the other hand, by the previous argument, if
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\neq\mathbb{S}^{n-1},$
we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\beta}}[\theta]:\theta\in\gamma_{\beta}\cup-\gamma_{\beta}\},$
hence, for any $\alpha\prec\beta$,
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}\subseteq\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\}$
$\subseteq$ $\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta}\cup-\overline{C}_{\beta}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\beta}}[\theta]:\theta\in\gamma_{\beta}\cup-\gamma_{\beta}\}.$
We now select a totally ordered subset of $\mathcal{G}$, which we denote by $\mathcal{H}$. Define
$\mathcal{F}:=\Bigl\{\alpha\in\mathcal{H}:\ \bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha},\ t\in[m,1]\}=\mathbb{S}^{n-1}\Bigr\}.$
We aim to show this set coincides with $\mathcal{H}$, i.e. $\mathcal{F}=\mathcal{H}$. Otherwise, there exists some $\beta_{0}\in\mathcal{H}$ such that for every $\alpha\prec\beta_{0}$ we have $\alpha\notin\mathcal{F}$, while for every $\beta\succ\beta_{0}$ we have $\beta\in\mathcal{F}$. However, for every $\alpha\prec\beta_{0}$ we have
$\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}\subseteq\bigcup\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\},$
and therefore
$\bigcup\{\Lambda_{t}[\theta]:\theta\in C_{\beta_{0}}\cup-C_{\beta_{0}}\ \text{and}\ t\in[m,1]\}$
$\subseteq$ $\bigcup_{\alpha\prec\beta_{0}}\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha}\cup-\overline{C}_{\alpha}\ \text{and}\ t\in[m,1]\}$
$\subseteq$ $\bigcup_{\alpha\prec\beta_{0}}\{\Lambda_{t_{\alpha}}[\theta]:\theta\in\gamma_{\alpha}\cup-\gamma_{\alpha}\}$
$\subseteq$ $\bigcup\{\Lambda_{t_{\beta_{0}}}[\theta]:\theta\in\gamma_{\beta_{0}}\cup-\gamma_{\beta_{0}}\},$
which yields
$\mathbb{S}^{n-1}$ $=\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\beta_{0}}\cup-\overline{C}_{\beta_{0}}\ \text{and}\ t\in[m,1]\}$
$\subseteq\bigcup\{\Lambda_{t_{\beta_{0}}}[\theta]:\theta\in\gamma_{\beta_{0}}\cup-\gamma_{\beta_{0}}\}\subseteq\Lambda_{\beta_{0}},$
contradicting the assumption.
Now define $\alpha_{0}$ as the infimum of $(\mathcal{H},\prec)$. Clearly, $C_{\alpha_{0}}=\emptyset$, and $\gamma_{\alpha_{0}}$ then satisfies
$\mathbb{S}^{n-1}$ $=\bigcup\{\Lambda_{t}[\theta]:\theta\in\overline{C}_{\alpha_{0}}\cup-\overline{C}_{\alpha_{0}}\ \text{and}\ t\in[m,1]\}$
$=\bigcup\{\Lambda_{t_{\alpha_{0}}}[\theta]:\theta\in\gamma_{\alpha_{0}}\cup-\gamma_{\alpha_{0}}\}\subseteq\Lambda_{\alpha_{0}},$
which contradicts our initial assumption.
Consequently, our assumption does not hold, and therefore Proposition (3) is valid. ∎
@@
$=$ $\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}$
@@
$=$ $\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c}$
@@
$=$ $t_{0}K\bigcup\left\{\displaystyle\bigcup_{\xi\in\mathbb{S}^{n-1}}\big[\xi^{\perp}\backslash\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\backslash t_{0}(K\cap\xi_{0}^{\perp})\big)\big]\right\}^{c},$